### v1

# iTBLS: A Dataset of Interactive Conversations Over Tabular Information

Abstract
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations situated in tables from scientific articles. This dataset is designed to facilitate human-AI collaborative problem-solving through AI-powered multi-task tabular capabilities. In contrast to prior work that models interactions as factoid QA or procedure synthesis, iTBLS broadens the scope of interactions to include mathematical reasoning, natural language manipulation, and expansion of existing tables from natural language conversation by delineating interactions into one of three tasks: interpretation, modification, or generation. Additionally, the paper presents a suite of baseline approaches to iTBLS, utilizing zero-shot prompting and parameter-efficient fine-tuning for different computing situations. We also introduce a novel multi-step approach and show how it can be leveraged in conjunction with parameter-efficient fine-tuning to achieve the state-of-the-art on iTBLS; outperforming standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
1 Introduction
Recent research on Conversational AI has focused on adding enhanced multi-task capabilities to large language models (LLMs). This research includes building systems capable of situated interactions over structured knowledge sources such as tabular information Sundar and Heck (2022). Automated methods for tabular interpretation, manipulation, and generation empower users by saving time and reducing errors in managing tabular content Kardas et al. (2020). Previous studies have focused on individual aspects of tabular data management: representation learning for interpretation tasks like grounded question answering, manipulation for data wrangling, and generation for summarizing textual information independently Nakamura et al. (2022a); Sundar and Heck (2023); Sundar et al. (2024); Fang et al. (2024).
The development of situated conversational interactions over tables necessitates a suite of approaches to unify tabular interpretation, modification, and generation in a conversational context. Additionally, an important yet largely unaddressed challenge in interacting with tabular sources is the ability to modify existing tabular content using conversational natural language commands.
To address these challenges, this paper introduces Interactive Tables (iTBLS)  [^1] The dataset and baseline models will be released with the camera-ready version under Apache License 2.0, a dataset of interactive conversations in English situated in tabular information. iTBLS decomposes the challenge into three distinct tasks: interpretation, which involves understanding tabular content within a conversational framework; modification, which entails manipulating tabular content through natural language commands; and generation, which focuses on integrating new natural language information into existing tables. The tabular information in iTBLS is sourced from scientific articles hosted on arXiv  [^2] https://arxiv.org, an open-access repository of academic preprints.
In contrast to earlier methods that approached tabular interpretation as question-answering focused on single-cell factoids, iTBLS introduces a more complex conversational interaction framework. Beyond factoid question-answering, iTBLS encompasses tasks such as comparison, determining absolute and relative positions, and mathematical reasoning. Previous research primarily examined procedural command generation for spreadsheets or the alignment of tabular data through LLMs. iTBLS integrates these functionalities into a unified task, enabling the manipulation of existing tables through natural-language commands. On tabular generation, while prior work addressed the summarization of natural language paragraphs in a tabular format, iTBLS focuses on generating row or column data conversationally.
Furthermore, this paper presents baseline approaches utilizing zero-shot prompting of off-the-shelf LLMs and parameter-efficient fine-tuning and represents a suite of techniques for different computing situations. Additionally, a novel multi-step approach is presented to improve accuracy and reduce errors, inspired by past work using multi-step methods to improve a variety of approaches including complex reasoning Jiang et al. (2023b), external tool use Paranjape et al. (2023), feedback Richardson et al. (2023), self-refinement Madaan et al. (2024), and mathematical reasoning Zhou et al. (2023). This multi-step method results in an improvement over standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
The contributions of this work are as follows:
  • Creating iTBLS, a dataset of tabular interactions unifying interpretation, modification, and generation
  • Extending prior tabular datasets by collecting information from arXiv
  • Broadening the scope of interactions to include mathematical reasoning, natural language manipulation, and natural language expansion
  • Completing a series of baselines utilizing zero-shot prompting, parameter-efficient fine-tuning, and a task-specific multi-step approach
The rest of the paper is organized as follows: Section 2 reviews prior research addressing tabular interpretation, modification, and generation. Section 3 presents a description of the tasks in iTBLS and outlines the dataset collection process. Sections 4 and 5 detail the baseline approaches to address iTBLS and their associated results. Lastly, Section 6 concludes the paper.
Figure 1: Examples of interactions from the Interactive Tables (iTBLS) dataset.
2 Related Work
A detailed survey of LLMs for tabular data is available in Fang et al. (2024). Related work on paired natural-language and tabular data can be broadly classified by the nature of the interaction: tabular interpretation, tabular modification, and tabular generation.
2.1 Tabular Interpretation
Tabular interpretation involves a dialogue turn focused on extracting information from a specific cell in a table, such as identifying a cell satisfying certain criteria. Prior research on tabular interpretation focused on grounded question-answering. An important challenge in the collection of such datasets is the availability of large-scale tabular data. Consequently, many tabular datasets are constructed from online resources such as Wikipedia including WikiTableQuestions Pasupat and Liang (2015), ManyModalQA Hannan et al. (2020), TaBERT Yin et al. (2020), NQ-Tables Herzig et al. (2021), FEVEROUS Aly et al. (2021), FeTaQA Nan et al. (2022), HybriDialogue Nakamura et al. (2022b), and HiTab Cheng et al. (2022). Other tabular datasets are constructed from financial reports including TAT-QA Zhu et al. (2021), FinQA Chen et al. (2021), MultiHiertt Zhao et al. (2022).
Proposed approaches to address the tabular interpretation task include architectures based off of the Transformer encoder Yin et al. (2020); Herzig et al. (2020); Chen et al. (2020); Eisenschlos et al. (2020); Liu et al. (2021); Gu et al. (2022); Yang et al. (2022), decoder Gong et al. (2020); Akhtar et al. (2023); Zha et al. (2023); Jiang et al. (2023a); Zhang et al. (2023b); Sui et al. (2024), or both (encoder-decoder) Nakamura et al. (2022b); Deng et al. (2022); Sundar and Heck (2023); Sarkar and Lausen (2023).
2.2 Tabular Modification
Tabular modification concerns the manipulation of the content within an existing table without altering the overall structure of rows and columns. Early work on tabular modification explored the generation of procedural commands for spreadsheets using synthesis algorithms Singh and Gulwani (2012); Shigarov and Mikhailov (2017). Tools utilizing programming-by-example to parse user intents into executable commands have also been explored Scaffidi et al. (2009); Kandel et al. (2011); Jin et al. (2017); Petricek et al. (2023); Chen et al. (2023). More recent work has shifted focus towards leveraging LLMs to synthesize commands for tools Huang et al. (2023), reformat tabular information Nobari and Rafiei (2023), and execute programming commands Liu et al. (2023).
2.3 Tabular Generation
Tabular generation focuses on expanding an existing table by adding a new row or column. Research on tabular generation initially employed discriminative techniques, such as tree-based methods for generating tables of contents Branavan et al. (2007) and SVMs to classify text across various labels Aramaki et al. (2009). Recent approaches have shifted towards neural techniques including Generative Adversarial Networks (GANs) Xu and Veeramachaneni (2018); Park et al. (2018); Chen et al. (2019); Zhao et al. (2021), Autoencoders Li et al. (2019); Darabi and Elor (2021), Diffusion models Kotelnikov et al. (2023), and LLMs Borisov et al. (2023); Solatorio and Dupriez (2023); Zhang et al. (2023a); Gulati and Roysdon (2023); Zhao et al. (2023); Seedat et al. (2024).
A similar line of research also explores the generation of tabular data from associated textual information. Wu et al. (2022) introduced four datasets and proposed a modification to the Transformer’s attention mechanism to summarize textual information in a tabular format by inverting datasets created for the dual task of converting tables to text, (as opposed to new conversational evidence). Other approaches to summarize textual information in a tabular format include the addition of learnable bias parameters Pietruszka et al. (2022), structure-aware instruction-tuning Tang et al. (2023), and reformulating the task within a question-answering framework Sundar et al. (2024).
In contrast to prior work addressing a single mode of interaction, iTBLS is a dataset unifying tabular interpretation, modification, and generation in a conversational format. Additionally, iTBLS broadens the range of interactions to include mathematical reasoning, natural language manipulation, and the expansion of tables using natural language. Furthermore, by leveraging scientific articles from arXiv as a primary source, iTBLS introduces a novel and rich source of information that is not present in existing datasets.
3 The iTBLS Dataset
The Interactive Tables (iTBLS) dataset features 3-turn conversational interactions situated in tabular data, covering the three distinct types of interactions described in Section2: interpretation, modification, and generation. Each example type is exemplified in Figure 1 and described below. In addition, since the mode of interaction is not known a priori, any proposed approach using iTBLS must effectively identify the interaction type, either explicitly or implicitly. In the following sections, we provide a detailed description of each type of interaction and outline the dataset collection process.
3.1 Tasks
Tabular Interpretation: In iTBLS, interpretive interactions are structured as question-answer pairs, where the goal is to identify the cell referred to by the question. The references could be absolute (referring to a specific row or column), or relative (referring to one cell in the context of another). Appendix A.1 details absolute and relative references in iTBLS.
Tabular modification: We conceptualize modification in iTBLS as a series of cell swaps, positing that any content rearrangement can ultimately be reduced to such exchanges. This approach allows for both explicit references, where specific row and column numbers are cited, and implicit references, which rely on the content or relative positions of cells. Table 7 in Appendix A.1 showcases examples from iTBLS. As observed, there is a mix of explicit and implicit references to the specific contents to be manipulated.
Tabular generation: In iTBLS, table generation is guided by new natural language evidence. This evidence clarifies appending a row or column, defines the suitable header, and supplies the data entries for the new row (or column) relative to existing columns (or rows). This process ensures that the added elements are contextually relevant and accurately integrated into the table. Table 8 in Appendix A.1 provides examples of such interactions, demonstrating how users can request the incorporation of new row and column data into an established table framework.
In iTBLS, the mode of interaction is not explicitly stated by the user, introducing an additional task: interaction identification. This task involves predicting whether the interaction is intended for interpretation, modification, or generation based solely on the user’s request.
3.2 Dataset Collection
To collect the dataset, first we use AxCell Kardas et al. (2020) an automatic machine learning pipeline for extracting results from papers. AxCell is used to parse tabular information from papers on arXiv to populate online leaderboards comparing scientific methods. Using AxCell, we collect 20,000 tables from academic papers in Mathematics, Physics, and Computer Science over a period spanning from 2007 to 2014. The tables are processed to remove stray characters resulting from the conversion from . Additionally, only tables with at least three rows and three columns to at most ten rows or ten columns are retained. The final dataset consists of 4000 tables split between train, development, and test sets.
For each table, we generate three sequential edits corresponding to different types of interaction. Interpretation involves generating a dialogue turn (question-answer pair) grounded on a single cell of the table. Modification involves manipulating two cells of an existing table by swapping them. Finally, generation encompasses the task of appending either a new row or a column to an existing table based on a natural language utterance.
To enhance the quality of the dataset and minimize errors, we implement a strategic selection process for the table components involved in each interaction. In interpretation, a cell is randomly selected to ground the dialogue. For modification, two cells are chosen and their positions are swapped to simulate a realistic table manipulation scenario. In generation, all cells in a randomly masked row or column are used as the basis for appending new table data. All of the interactions are based on cells that do not belong to row or column headers, that is, they reside in the body of the table.
For our dataset creation, we employ two distinct sources for generating dialogue turns based on the type of interaction and the specific table component involved. For tasks related to tabular interpretation and modification, we engage crowd-workers from Amazon Mechanical Turk (AMT). These workers are tasked with formulating questions or commands that pertain to the pre-identified cell(s) designated for each interaction. We recruit workers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the USA. Each crowdworker is compensated at a rate of \$0.15 per Human Intelligence Task (HIT), with the average completion time for each HIT being approximately 40 seconds. Detailed information on the AMT interface used for these tasks is included in Appendix A.3.
For generation, GPT-4 is prompted to write a dialogue turn summarizing a row or column of the table. We prompt the version of GPT-4 as of 2/16/2024 using the OpenAI API. The prompt is as follows:
The string contains information from a table [table]. Describe the content in this [row/column] for a visually impaired user in one line. Make sure to include all information from the rows and columns and appropriate headers so the user can understand the content.
Each sample in the dataset contains the source arXiv ID, the table that the conversation is situated in, the index of that table within the paper (e.g. Table X), the utterance describing the interaction, the ground truth cell(s) involved in the interaction, and finally the expected output. Statistics of the dataset are provided in Table 1.
Interpret Modify Generate
train dev test train dev test train dev test
# Samples 3001 646 521 3001 646 521 3001 646 521
# Per Utterance
Tokens 14.3 14.3 14.3 18.5 17.7 17.8 59.3 60.6 55.9
Words 10.7 10.7 10.3 13.5 13.2 13.0 32.0 31.8 29.5
# Per Input Table
Cells 28.0 28.8 27.7 28.0 28.8 27.5 25.2 26.0 25.1
Columns 5.0 5.0 4.9 5.0 5.0 4.9 4.8 4.8 4.7
Rows 5.5 5.5 5.6 5.5 5.5 5.6 5.3 5.3 5.3
Table 1: Dataset Statistics
Figure 2: iTBLS multi-step approach. The input table and user request are first converted into an encoded form by an LLM. The encoded request is then used to generate the output via a task-specific procedure in python.
4 Baseline Approaches
[…]
4.1 Zero-shot prompting
Zero-shot prompting leverages off-the-shelf LLMs that have undergone pre-training with a causal language modeling objective followed by supervised fine-tuning to align with human instructions. The LLM is not fine-tuned on iTBLS and is directly prompted without examples in-context. The prompt consists of the input table, the interaction request, and additional task-specific instructions to ensure the output is formatted correctly.
4.2 Parameter-efficient fine-tuning (PEFT)
We use QLoRA Dettmers et al. (2023), a parameter-efficient approach to fine-tune LLMs. In this method, an LLM is fine-tuned for each sub-task by maximizing a causal language modeling objective based on the specific interaction. The objective to be optimized is described in Section 5.
4.3 Multi-step approach
The Multi-Step approach (MS) breaks down the entire task into a series of intermediate steps, addressing part of the specific task. For interpretation and modification, this process first identifies the cells referred to by the user using an LLM in a zero-shot configuration followed by program synthesis to execute the action. However, due to the custom encoding used in identifying cell references (Figure 2), the approach did not yield satisfactory results when compared to zero-shot prompting. For generation, the multi-step approach breaks down table generation as two-step question-answering. Additional details are provided in Section 5.3.
4.4 Multi-Step + PEFT
This approach first breaks down the entire process into a sequence of intermediate steps. Next, an LLM is fine-tuned using parameter-efficient fine-tuning to address each step independently.
5 Results
We introduce a modular approach where we use an LLM to identify the nature of the interaction followed by another LLM to execute the operation described in the user utterance. A high-level overview is depicted in Figure 3. We use the version of Llama 2 Touvron et al. (2023) already pre-trained and instruction-aligned using supervised fine-tuning and reinforcement learning with human feedback for dialogue use cases (Llama-2-[7/13/70]B-chat-hf) [^3] https://huggingface.co/meta-llama, hereafter referred to as Llama-2-[7/13/70]B.
Figure 3: A block diagram of our approach to iTBLS
5.1 Interaction Identification
For interaction identification we benchmark both zero-shot and fine-tuned approaches using GPT-{3.5,4} and Llama 2-{7B,13B,70B}. We use classification accuracy as the metric. For zero-shot experiments, we instruct the model to wrap the identified task in special tags, e.g. “[TASK] modify [/TASK]". This enables automated extraction of the answers in the zero-shot setting. Any response that fails to follow the requested format is considered incorrect. We aggregate all data across tasks into pairs of requests $i$ and tasks $k$ $(r_{i,k},t_{k})$ and maximize the objective:
$p(t_{k}|r_{i,k})$ (1)
Results for interaction identification are provided in Table 2 where PEFT using Llama 2 achieves near-perfect performance on the task.
Model Meth. Split Accuracy
Llama-2-7b ZS dev 0.531
test 0.516
Llama-2-13b ZS dev 0.163
test 0.150
Llama-2-70b ZS dev 0.501
test 0.483
GPT-3.5 turbo ZS dev 0.431
test 0.416
GPT-4 ZS dev 0.658
test 0.658
Llama-2-7b PEFT dev 0.997
test 0.997
Table 2: Interaction ID: Zero-shot (ZS) vs PEFT
5.2 Tabular Interpretation
Table 3 contains results for the zero-shot approach using the different sizes of Llama 2. For each model, we report the exact match score that identifies whether the entire answer string is present in a continuous form in the model’s response. Using larger models results in better exact match scores, with the answer phrase present in almost 30% of the responses from the 70B variant of Llama 2. We also report results using PEFT. In this situation, Llama 2 is fine-tuned to maximize the probability of the correct answer over the set of all possible answers $\mathcal{A}$ given a table and a user request:
$\operatorname*{argmax}_{\mathcal{A}}p(\text{answer}|\text{table},\text{request})$ (2)
Hyperparameters for fine-tuning are detailed in Appendix A.2. As expected, fine-tuning using QLoRA performs better than zero-shot prompting, achieving up to 2x the exact match score when compared to the zero-shot approaches (Table 3). Our best-performing method is a multi-step approach combined with parameter-efficient fine-tuning. In this method, we identify the cell referred to by the user utterance in the concept of the table (row and column number) and then generate a response based on this cell. Performing this operation ensures that the output response is necessarily a cell of the table as opposed to free-form generation with no constraint on the emitted LLM response being a cell of the input table. An example of the codified representation is available in Figure 2.
Model Method Split Exact
Match
Llama-2-7b ZS dev 0.176
test 0.199
Llama-2-13b ZS dev 0.170
test 0.213
Llama-2-70b ZS dev 0.233
test 0.291
Llama-2-7b PEFT dev 0.517
test 0.560
Llama-2-7b MS+PEFT dev 0.667
test 0.631
Table 3: Results: iTBLS-interpretation
5.3 Tabular Generation
First, we experiment with zero-shot prompting with Llama 2 to incorporate the row or column information into the table in a single step. During evaluation, we assume tabular information is order invariant, provided the contents align accurately with the correct header cells. To assess the effectiveness of our approach, we report the F1 score, that is, the harmonic mean of precision and recall computed between the cells of the predicted row or column and the ground truth. As evidenced by the results in Table 5, this single-step zero-shot prompting results in tables that are malformed, that is, the generated tables are inconsistent and incorrectly formatted, or do not match the dimensions of the expected ground truth to be generated. The malformed rate reduces as we experiment with larger models, going down from 88% to $\approx$ 26% with a marginal improvement in F1 scores.
To address this issue, we experiment with a multi-step approach. Prior work has shown the effectiveness of reformulating the table-generation task as two-stage machine reading comprehension Sundar et al. (2024) to achieve tables that are always well-formed. The first stage identifies whether the input utterance asks to generate a row or column using Llama 2 in a zero-shot configuration. Then, based on whether a row or column is to be generated, questions are synthesized using the header rows of the corresponding column or row. The synthesized questions are then answered using Llama 2 in a zero-shot configuration to yield the final output. Since we constrain the generated content to conform to the existing tabular structure, none of the tables contain formatting errors, resulting in a malformed rate of 0%.
To further improve results, we experiment with PEFT using QLoRA. The fine-tuning process is performed to ensure the LLMs emit cells of a table autoregressively by optimizing the objective in Equation 3. The output is parsed to reformat the CSV-serialization back into the original two-dimensional representation.
$\operatorname*{argmax}_{h}p(\text{cell}_{n}|\{\text{cell}_{ij}\},\text{table},\text{request})$ (3)
This results in a 2x improvement in F1 scores over the multi-step approach (Table 5), but at the cost of increased malformed rates. To alleviate this issue, we combine the multi-step and fine-tuning approaches. First, Llama 2 is used in a zero-shot configuration to determine whether a row or column is to be added to the table. Next, the question-answering reformulation is used to synthesize question-answer pairs. Llama 2 is then fine-tuned using QLoRA to answer these questions. This approach marries the best of both worlds, resulting in a further 2x improvement in F1 scores over the end-to-end fine-tuning while bringing the malformed rates to 0%.
5.4 Tabular Modification
Model Method Split Acc. Malf. Rate(%)
Llama-2-7b ZS dev 0.006 64.7
test 0.000 14.0
Llama-2-13b ZS dev 0.014 44.1
test 0.002 12.9
Llama-2-70b ZS dev 0.037 15.6
test 0.031 8.1
Llama-2-7b PEFT dev 0.139 2.5
test 0.140 3.1
Llama-2-7b MS+ PEFT dev 0.274 0
test 0.328 0
Table 4: Results: iTBLS-modification
For the modification task, we experiment with a fine-tuning approach with Llama-2-7b, as well as zero-shot approaches with three llama-2 models (7b, 13b, and 70b). For each experiment, we compute table modification accuracy. An output table is accurate when all row and column headers match the ground truth, as well as all cell values within the table. This metric is computed by converting the textual form of the output table into a DataFrame using the pandas library in python and then comparing it to ground truth. If the conversion to DataFrame fails, the table is considered malformed.
The accuracy and malformed rate are reported in Table 4. Zero-shot approaches with all three models achieve low (<4%) accuracy on the modification task, including a 0% accuracy from the smallest model (7B). This is not unexpected due to the need to identify cells from relative references in natural language text form and perform the correct swap, combined with the strict constraint of generating well-formed tables. All this must be performed while maintaining the remainder of the table’s values. This is further exacerbated by the presence of malformed table outputs, which make up as much as 15.6% of output tables in some experiments. Table 4 further demonstrates the strength of PEFT, where the standard approach boosts accuracy by 10.9% on the test split while using 1/10 the parameters. However, malformed tables remain a problem even with standard PEFT. This is addressed with a multi-step approach. First, we generate an encoding of the requested modification. This is then followed by a program synthesis to perform the swap on the table and return the resulting modified table. An example of the encoding to swap contents in the 3rd and 4th columns of the first row is: modify<SEP>1<SEP>4<SEP>1<SEP>3. The corresponding program synthesis is executed in pandas using df.iloc [^4] https://pandas.pydata.org/pandas-docs/stable/reference/api/pandas.DataFrame.iloc.html.
With our multi-step approach, we are able to achieve a 0% malformed rate (this is guaranteed by the method), and an 18.8% absolute increase in accuracy on the test split, thus demonstrating the value of the multi-step approach for the task of tabular modification.
Model Meth. Split F1 Malf. Rate(%)
Llama-2-7b ZS dev 0.000 87.9
test 0.000 87.9
Llama-2-13b ZS dev 0.003 57.8
test 0.005 62.1
Llama-2-70b ZS dev 0.063 29.5
test 0.078 22.6
Llama-2-7b MS dev 0.163 0.0
test 0.128 0.0
Llama-2-13b MS dev 0.094 0.0
test 0.073 0.0
Llama-2-70b MS dev 0.177 0.0
test 0.129 0.0
Llama-2-7b PEFT dev 0.309 12.0
test 0.315 4.9
Llama-2-7b MS+ PEFT dev 0.682 0.0
test 0.699 0.0
Table 5: Comparison of approaches for iTBLS-generation, ZS=Zero-shot, MS=Multi-step, PEFT=Parameter-efficient Fine-tuning
6 Conclusion
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations situated over tables in scientific papers. iTBLS consists of multi-turn conversations with three modes of interaction - interpretation, modification, and generation. iTBLS extends prior work by introducing new interactions such as comparison of table contents, mathematical and positional reasoning, manipulation using natural language commands, and generation of content from new evidence. Additionally, a set of baseline approaches for a variety of computing situations addressing iTBLS is presented. The baseline approaches utilize zero-shot prompting and parameter-efficient fine-tuning. Furthermore, a novel multi-step task-specific approach is introduced. Leveraging the multi-step process in conjunction with parameter-efficient fine-tuning yields the best performance on iTBLS, outperforming standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
7 Limitations
While iTBLS introduces a dataset for interactive conversations over tabular information, there are some avenues for improvement. In this dataset, modification is modeled as a series of swaps. A more comprehensive sequence of manipulations includes in-place modification of values and modifying a cell’s value based on other cells using both absolute and relative references. While sourcing tabular information from arXiv provides a cost-efficient approach, LLMs are often pre-trained on sources from arXiv. This paper alleviates the issue by sourcing natural language commands from crowdworkers. Future work could look at collecting tabular information from crowdworkers as well. While we present a suite of baseline approaches for iTBLS, there is still headroom between the presented approaches and perfect performance. We identify the closure of this gap as an avenue for future work.
Appendix A Appendix
We acknowledge the use of GitHub Copilot to assist in code completion.
A.1 Dataset Examples
Example
1 What is the 2nd cell value for row 4?
2 Tell me the final value in the column labeled k
3 What is the value of the cell to the left of the cell in the bottom right of the table.
Table 6: Example interactions in the Interpretation task of iTBLS
Example
1 The rows 1 and 4 in the Column “Citation" were accidentally switched. Please rectify the positions of these values so they are where they need to be.
2 Swap the contents of the second and last cell under repetitions.
3 Two values in the MCBLp column were put in the reverse spots. I need the values for the FM and PCC rows flipped.
Table 7: Example interactions in the Modification task of iTBLS
Example
1 The row 3 of the table shows the values for Peak as 4, X coordinate as 0.100, Y coordinate as -0.150, A as 0.5, standard deviation ($\sigma$) as 0.02, and Local lnZ as -7.824.
2 The column “Method 2 (with sub-clustering)" contains the ‘Nlike’ values in different rows: 27,658 in the second row, 69,094 in the third row, 579,208 in the fourth row, and 43,093,230 in the fifth row, while the remaining rows from six to nine contain no data (NaN).
3 The column R contains eight numerical values in increasing order: 3.34, 3.40, 3.66, 5.06, 6.02, 6.61, 4.05, and 4.11.
Table 8: Example interactions in the Generation task of iTBLS
A.2 Hyperparameters
Parameter Efficient Fine-tuning
Parameter Value
Rank 64
$\alpha$ 16
Dropout 0.1
Table 9: Hyperparameters for Parameter-efficient fine-tuning using QLoRA
Additional hyperparameters:
Parameter Value
Learning Rate 2e-4
Batch size 4
Warmup Schedule Constant
Warmup Ratio 0.03
Epochs 5
Optimizer paged_adamw_32bit [^5] https://huggingface.co/docs/bitsandbytes/main/en/reference/optim/adamw
Compute 8 Nvidia A40 GPUs
Table 10: Additional hyperparameters for fine-tuning experiments
A.3 Mechanical Turk Interface
Figure 4: Amazon Mechanical Turk Interface to collect iTBLS interpretation
Figure 5: Amazon Mechanical Turk Interface to collect iTBLS modification