### v2

# iTBLS: A Dataset of Interactive Conversations Over Tabular Information

Abstract
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations that focuses on natural-language manipulation of tabular information sourced from academic pre-prints on ArXiv. The iTBLS dataset consists of three types of tabular tasks – interpretation, modification, and generation. Interpretation focuses on tabular understanding, modification focuses on manipulating tabular information, and generation focuses on the addition of new natural-language evidence. In addition, the paper presents a novel framework that reformulates tabular operations as question-answering, where an appropriate question is formulated based on the nature of interaction and the question is answered using the user request as evidence. The developed approach results in an improvement on all tasks on a sequence-to-sequence modeling baseline on iTBLS. In addition, the question-answering-based reformulation is applied to datasets from prior work for the text-to-table task where textual paragraphs are summarized into tables. The novel approach results in up to 13% improvement in Exact-Match accuracy and up to 16% improvement in BERTScores compared to the prior state-of-the-art.
iTBLS: A Dataset of Interactive Conversations Over Tabular Information
Anirudh Sundar1  and Christopher Richardson2 thanks: Work done while at Georgia Tech  and Adar Avsian1  and Larry Heck1 1 Georgia Institute of Technology, USA 2 Google Inc., USA asundar34, larryheck@gatech.edu
1 Introduction
Recent research on Conversational AI has focused on adding enhanced multi-task capabilities to large language models (LLMs). This research includes building systems capable of situated interactions over structured knowledge sources such as tabular information Sundar and Heck (2022). Automated methods for tabular interpretation, manipulation, and generation empower users by saving time and reducing errors in managing tabular content Kardas et al. (2020). Previous studies have focused on individual aspects of tabular data management: representation learning for interpretation tasks like grounded question answering, manipulation for data wrangling, and generation for summarizing textual information independently Nakamura et al. (2022a); Sundar and Heck (2023); Fang et al. (2024).
The development of situated conversational interactions over tables necessitates a suite of approaches to unify tabular interpretation, modification, and generation in a conversational context. Additionally, an important yet largely unaddressed challenge in interacting with tabular sources is the ability to modify existing tabular content using conversational natural language commands.
To address these challenges, this paper introduces Interactive Tables (iTBLS)  [^1] https://huggingface.co/datasets/avalab/iTBLS, a dataset of interactive conversations in English situated in tabular information. iTBLS decomposes the challenge into three distinct tasks: interpretation, which involves understanding tabular content within a conversational framework; modification, which entails manipulating tabular content through natural language commands; and generation, which focuses on integrating new natural language information into existing tables. The tabular information in iTBLS is sourced from scientific articles hosted on arXiv  [^2] https://arxiv.org, an open-access repository of academic preprints.
Beyond factoid question-answering, iTBLS encompasses tasks such as comparison, determining absolute and relative positions, and mathematical reasoning. Previous research primarily examined procedural command generation for spreadsheets or the alignment of tabular data through LLMs. iTBLS integrates these functionalities into a unified task, enabling the manipulation of existing tables through natural-language commands. On tabular generation, while prior work addressed the summarization of natural language paragraphs in a tabular format, iTBLS focuses on generating row or column data conversationally.
In addition to building iTBLS, this paper develops a novel approach to address tabular operations by reformulating the task as conditional question answering. Furthermore, the question-answering-based reformulation is applied to other datasets introduced in prior work Wu et al. (2022) and results in better performance in terms of both table-cell accuracy and BERTScore.
Figure 1: Examples of interactions from the Interactive Tables (iTBLS) dataset.
The contributions of this work are as follows:
  • Creating iTBLS, a dataset of tabular interactions unifying interpretation, modification, and generation.
  • Extending prior tabular datasets by collecting information from arXiv
  • Broadening the scope of interactions to include mathematical reasoning, natural language manipulation, and natural language expansion.
  • Introducing a novel approach for table generation tasks through a two-stage reformulation that first identifies the cells to be manipulated and generates a question based on the requested operation, then answers those questions using the user request and the input table as evidence.
  • Demonstrating up to 13% improvement in table-cell accuracy and up to 16% improvement in BERTScore using the novel approach on the text-to-table task introduced by prior work.
2 Related Work
A detailed survey of LLMs for tabular data is available in Fang et al. (2024). Related work on paired natural-language and tabular data can be broadly classified by the nature of the interaction: tabular interpretation, tabular modification, and tabular generation.
2.1 Tabular Interpretation
Tabular interpretation involves a dialogue turn focused on extracting information from a specific cell in a table, such as identifying a cell satisfying certain criteria. Prior research on tabular interpretation focused on grounded question-answering. An important challenge in the collection of such datasets is the availability of large-scale tabular data. Consequently, many tabular datasets are constructed from online resources such as Wikipedia including WikiTableQuestions Pasupat and Liang (2015), ManyModalQA Hannan et al. (2020), TaBERT Yin et al. (2020), NQ-Tables Herzig et al. (2021), FEVEROUS Aly et al. (2021), FeTaQA Nan et al. (2022), HybriDialogue Nakamura et al. (2022b), and HiTab Cheng et al. (2022). Other tabular datasets are constructed from financial reports including TAT-QA Zhu et al. (2021), FinQA Chen et al. (2021), MultiHiertt Zhao et al. (2022), or scientific reviews Sundar et al. (2024).
Proposed approaches to address the tabular interpretation task include architectures based off of the Transformer encoder Yin et al. (2020); Herzig et al. (2020); Chen et al. (2019b); Eisenschlos et al. (2020); Liu et al. (2021); Gu et al. (2022); Yang et al. (2022), decoder Gong et al. (2020); Akhtar et al. (2023); Zha et al. (2023); Jiang et al. (2023); Zhang et al. (2023); Sui et al. (2024); Cremaschi et al. (2025), or both (encoder-decoder) Nakamura et al. (2022b); Deng et al. (2022); Sundar and Heck (2023).
2.2 Tabular Modification
Tabular modification concerns the manipulation of the content within an existing table without altering the overall structure of rows and columns. Early work on tabular modification explored the generation of procedural commands for spreadsheets using synthesis algorithms Singh and Gulwani (2012); Shigarov et al. (2019). Tools utilizing programming-by-example to parse user intents into executable commands have also been explored Scaffidi et al. (2009); Kandel et al. (2011); Jin et al. (2017); Petricek et al. (2023); Chen et al. (2023); Xing et al. (2024). More recent work has shifted focus towards leveraging LLMs to synthesize commands for tools Huang et al. (2024), reformat tabular information Dargahi Nobari and Rafiei (2024), and execute programming commands Liu et al. (2024).
2.3 Tabular Generation
Tabular generation focuses on expanding an existing table by adding a new row or column. Research on tabular generation initially employed discriminative techniques, such as tree-based methods for generating tables of contents Branavan et al. (2007) and SVMs to classify text across various labels Aramaki et al. (2009). Recent approaches have shifted towards neural techniques including Generative Adversarial Networks (GANs) Xu and Veeramachaneni (2018); Park et al. (2018); Chen et al. (2019a); Zhao et al. (2021), Autoencoders Li et al. (2019); Darabi and Elor (2021), Diffusion models Kotelnikov et al. (2023), and LLMs Borisov et al. (2023); Solatorio and Dupriez (2023); Gulati and Roysdon (2023); Zhao et al. (2023); Seedat et al. (2024); Deng et al. (2024).
A similar line of research also explores the generation of tabular data from associated textual information. Wu et al. (2022) introduced four datasets and proposed a modification to the Transformer’s attention mechanism to summarize textual information in a tabular format by inverting datasets created for the dual task of converting tables to text, (as opposed to new conversational evidence). Other approaches to summarize textual information in a tabular format include the addition of learnable bias parameters Pietruszka et al. (2022) and structure-aware instruction-tuning Tang et al. (2023).
In contrast to prior work addressing a single mode of interaction, iTBLS is a dataset unifying tabular interpretation, modification, and generation in a conversational format. Additionally, iTBLS broadens the range of interactions to include mathematical reasoning, natural language manipulation, and the expansion of tables using natural language. Furthermore, by leveraging scientific articles from arXiv as a primary source, iTBLS introduces a novel and rich source of information that is not present in existing datasets.
3 The iTBLS Dataset
The Interactive Tables (iTBLS) dataset features conversational interactions situated in tabular data, covering the three distinct types of interactions described in Section 2: interpretation, modification, and generation. Each example type is exemplified in Figure 1 and described below. In addition, since the mode of interaction is not known a priori, any proposed approach using iTBLS must effectively identify the interaction type, either explicitly or implicitly. In the following sections, we provide a detailed description of each type of interaction and outline the dataset collection process.
3.1 Tasks
Tabular Interpretation: In iTBLS, interpretive interactions are structured as question-answer pairs, where the goal is to identify the cell referred to by the question. The references could be absolute (referring to a specific row or column), or relative (referring to one cell in the context of another). Appendix LABEL:sec:dataset_examples_itbls details absolute and relative references in iTBLS.
Tabular modification: We conceptualize modification in iTBLS as a series of cell swaps, positing that any content rearrangement can ultimately be reduced to such exchanges. This approach allows for both explicit references, where specific row and column numbers are cited, and implicit references, which rely on the content or relative positions of cells. Table LABEL:tab:modif_example in Appendix LABEL:sec:dataset_examples_itbls showcases examples from iTBLS. As observed, there is a mix of explicit and implicit references to the specific contents to be manipulated.
Tabular generation: In iTBLS, table generation is guided by new natural language evidence. This evidence clarifies appending a row or column, defines the suitable header, and supplies the data entries for the new row (or column) relative to existing columns (or rows). This process ensures that the added elements are contextually relevant and accurately integrated into the table. Table LABEL:tab:generat_example in Appendix LABEL:sec:dataset_examples_itbls provides examples of such interactions, demonstrating how users can request the incorporation of new row and column data into an established table framework.
In iTBLS, the mode of interaction is not explicitly stated by the user, introducing an additional task: interaction identification. This task involves predicting whether the interaction is intended for interpretation, modification, or generation based solely on the user’s request.
3.2 Dataset Collection
To collect the dataset, first we use AxCell Kardas et al. (2020) an automatic machine learning pipeline for extracting results from papers. AxCell is used to parse tabular information from papers on arXiv to populate online leaderboards comparing scientific methods. Using AxCell, we collect 20,000 tables from academic papers in Mathematics, Physics, and Computer Science over a period spanning from 2007 to 2014. The tables are processed to remove stray characters resulting from the conversion from . Additionally, only tables with at least three rows and three columns to at most ten rows or ten columns are retained. The final dataset consists of 4000 tables split between train, development, and test sets.
For each table, we generate three sequential edits corresponding to different types of interaction. Interpretation involves generating a dialogue turn (question-answer pair) grounded on a single cell of the table. Modification involves manipulating two cells of an existing table by swapping them. Finally, generation encompasses the task of appending either a new row or a column to an existing table based on a natural language utterance.
To enhance the quality of the dataset and minimize errors, we implement a strategic selection process for the table components involved in each interaction. In interpretation, a cell is randomly selected to ground the dialogue. For modification, two cells are chosen and their positions are swapped to simulate a realistic table manipulation scenario. In generation, all cells in a randomly masked row or column are used as the basis for appending new table data. All of the interactions are based on cells that do not belong to row or column headers, that is, they reside in the body of the table.
For our dataset creation, we employ two distinct sources for generating dialogue turns based on the type of interaction and the specific table component involved. For tasks related to tabular interpretation and modification, we engage crowd-workers from Amazon Mechanical Turk (AMT). These workers are tasked with formulating questions or commands that pertain to the pre-identified cell(s) designated for each interaction. We recruit workers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the USA. Each crowdworker is compensated at a rate of \$0.15 per Human Intelligence Task (HIT), with the average completion time for each HIT being approximately 40 seconds. Detailed information on the AMT interface used for these tasks is included in Appendix LABEL:sec:app_mturk.
For generation, GPT-4 is prompted to write a dialogue turn summarizing a row or column of the table. The prompt is as follows:
The string contains information from a table [table]. Describe the content in this [row/column] for a visually impaired user in one line. Make sure to include all information from the rows and columns and appropriate headers so the user can understand the content.
Each sample in the dataset contains the source arXiv ID, the table that the conversation is situated in, the index of that table within the paper (e.g. Table X), the utterance describing the interaction, the ground truth cell(s) involved in the interaction, and finally the expected output. Statistics of the datasets are provided in Table 1.
Statistic Interpret Modify Generate
# Samples 4168 4168 4168
# Per utterance
Words 10.6 13.4 31.6
Tokens 14.3 18.3 59.1
# Per table
Cells 28.1 28.1 25.31
(Cols/Rows) 5.0/5.5 5.0/5.5 4.8/5.3
Table 1: Statistics of the iTBLS dataset
4 Methods
4.1 Table operations through conditional question answering
We also present a novel approach that reformulates operations on tables as question answering. A primary challenge in tabular operations using LLMs lies in ensuring the syntactic validity of the produced tables. Every row and column in a table must contain the same number of cells, with row and column headers delineating relationships between cells. Failing to adhere to this constraint invalidates the structure of the table and the information presented. Prior work addresses this constraint by including additional parameters like row and column relation embeddings Wu et al. (2022) or positional bias Pietruszka et al. (2022) to get the model to attend to header cells while generating content. However, this results in highly specialized architectures for a singular task. Breaking the task down into question-answering results in a more interpretable framework while ensuring validity of the generated tables.
The first step identifies the mode of interaction and the cell(s) the user is referring to, which is used to formulate a question. The second step converts the table into a pandas dataframe, parses the table and the question generated from the previous step to obtain a pandas command corresponding to the task, and executes the command on the dataframe to generate the final table. Generating a valid command ensures that the final table is syntactically valid as well (that is, the number of columns across all rows is consistent).
For the interpret task, the question-answer reformulation is trivial, since all interpretive queries and associated responses are naturally question-answer pairs. For the modify task, the question is of the form To which cells is the user referring?. A language model is then fine-tuned to generate a response containing the cells (indexed by row and column). Then, the LLM response is reformatted into an appropriate pandas command. Finally, for the generate task, the question-answering is more nuanced. First, the user request is parsed to identify whether a row or a column is to be appended. The header of the corresponding row is then extracted from the user request. Using the extracted header and the other header cells of the table, questions are generated for each of the empty cells to be filled in the form What is the row value for column?. The user request is parsed to obtain the answers to these generated questions, forming the corresponding row or column to be appended.
Figure 2: Overview of the novel question-answering reformulation to perform table operations
5 Results
5.1 Experimental Setup
For our experiments, we utilize Gemma models Team et al. (2024). We fine-tune the instruction-tuned base model gemma-2-9b-it using LoRA Hu et al. (2022). Hyperparameters for our training setup as well as LoRA parameters are shown in Appendix A.
5.2 Datasets
In addition to the iTBLS dataset, we also evaluate our method on five datasets to summarize textual paragraphs to tables Wu et al. (2022). While iTBLS is a table-to-table or table-to-text task, the datasets proposed by Wu et al. (2022) address the dual problem of text-to-table. The datasets consist of textual paragraphs containing some information that is to be converted into a tabular format by determining both the appropriate header cells and the content that the table is filled with.
Wu et al. (2022) present datasets for the text-to-table task by inverting datasets created for the dual problem of generating textual descriptions from tables. Each dataset consists of textual paragraphs paired with tabular information summarizing content in the text. Dataset statistics are available in Appendix A.3. Each dataset is described below.
E2E Novikova et al. (2017) concerns restaurant descriptions, requiring summarization of information into tables with descriptors like restaurant name, customer rating, and location. WikiTableText (WTT) Bao et al. (2018), sourced from Wikipedia, consists of natural language descriptions generated from tabular data across various topics. WikiBio Lebret et al. (2016) comprises introductions of individuals from Wikipedia alongside tabular summaries extracted from the same page’s information box. In contrast to E2E, the table headers in the WikiTableText and WikiBio datasets vary widely across data samples.
Example textual paragraphs and associated tables from each dataset are presented in A.4.
5.3 Metrics
Exact-Match (EM): On the iTBLS dataset, we report exact-match, that is, whether or not the generated table matches the ground-truth table exactly.
BERTScore: On the E2E, WTT, WikiBio and RotoWire datasets, we report BERTScore Zhang et al. (2020) in addition to EM to be consistent with prior work. BERTScore is a measure of semantic similarity which computes the similarity of embeddings in a latent space obtained using an encoder language model.
Consistent with prior work, all our evaluations are order-invariant. That is, credit is given as long as the generated cells are indexed by the correct row and column headers, even if the headers themselves are in different positions between the model-generated response and the ground-truth.
5.4 iTBLS
Results on the iTBLS dataset using a vanilla sequence-to-sequence approach and the question-answering-based method are presented in Table 2. As observed in the results, the generate task is the hardest, with performance slightly lower on the generate task when compared to interpret and modify. This is a result of the fact that the exact-match metric only provides credit when all cells are correct (necessitating that all cells in the output are identical to the ground truth) and does not provide partial credit for getting some of the cells right, and the fact that the generate task requires getting more cells right in comparison to the other tasks.
Split Approach Exact-Match
Interpret Seq2seq 88.29
iTBLS as QA 90.98
Modify Seq2seq 74.65
iTBLS as QA 89.58
Generate Seq2seq 48.94
iTBLS as QA 73.32
Table 2: Comparison between the question-answering reformulation and a vanilla sequence-to-sequence modeling approach on the iTBLS dataset
5.5 Text-to-table
The results on the text-to-table datasets proposed by Wu et al. (2022) are available in Table 3. Our method performs on par with or better than the prior state-of-the-art method in terms of BERTScore and is competitive with prior work in terms of Exact-Match. The exact-match score does not reflect true performance on the WikiBio dataset since synonyms are penalized under this framework. A deep-dive into the results is presented in Section 5.6.
Dataset Approach EM BS
WTT Wu et al. (2022) 62.71 80.74
Ours 75.96 95.52
Wikibio Wu et al. (2022) 69.71 76.56
Ours 66.65 92.60
E2E Wu et al. (2022) 97.94 98.57
Ours 97.64 99.35
Table 3: Comparison between our method and prior work on the text to table task in terms of Exact-Match and BERTScore
5.6 Analysis of Errors
An analysis of the difference in performance between the prior state of the art and our approach is presented in Table 6. As observed, the dataset is inconsistent in the description of individuals, with no consistent pattern when middle and last names are present. Furthermore, the use of quantifying information in the header as opposed to the table cell results in no credit using the exact-match metric, though the information contained is exactly the same between the prediction and the ground-truth. Finally, the datasets often contain textual examples with multiple possible tabular summarizations, all of which are equally valid, further complicating evaluation. In the third example in Table 6, the model correctly generates the ‘Occupation’ as a table header while the ground truth contains an erroneous sample, using the phrase ‘Known for’ instead of ‘Known as’.
Examples of errors in the iTBLS dataset are provided in Tables 4, 5, and LABEL:tab:difference_answers_itbls_generate. On the interpret task, the model incorrectly understand the user request, and produces the cell immediately to the right instead of three columns over. On the modify task (Table 5), the model incorrectly understands the references and swaps index (2,3) with (3,2) instead of swapping indices (2,2) and (3,3). On the generate task (Table LABEL:tab:difference_answers_itbls_generate), the model incorrectly places a tuple and hallucinates a value instead of performing the requested action.
Text: What is the value of the cell in row 1 that is three cells to the right of the cell with a value of 12%?
Input Table: row ID $\sigma\mu[I0]$ $\mu[\tau s]$ $\sigma[\tau s]$ $\sigma\mu[\tau s]$ 0 13% 912.5 $\mu s$ 91.9 $\mu s$ 10.1% 1 12% 18335.7 $\mu s$ 90.7 $\mu s$ 10.0% 2 12% 903.1 $\mu s$ 1832.7 $\mu s$ 10.0% Ground Truth: 10.0% Prediction: 18335.7 $\mu s$
Table 4: Example error for iTBLS interpret task. Table source: https://arxiv.org/pdf/1411.5458
Text: Swap row 1 in the second column with row 2 in the third column
Input Table: row ID col 1 col 2 col 3 0 X O X 1 NaN O O 2 O X X Ground Truth: row ID col 1 col 2 col 3 0 X O X 1 NaN X O 2 O X O Prediction: row ID col 1 col 2 col 3 0 X O X 1 NaN O X 2 O O X
Table 5: Example error for iTBLS modify task. Table source: https://arxiv.org/pdf/1411.4023
1. Text: Walter Clarence Henderson (28 February 1891 – 20 September 1968) was a progressive conservative party member of the Canadian house of commons. He was born in Carberry, Manitoba and became a farmer by career. He was elected at the Cariboo riding in the 1958 general election, defeating social credit incumbent Bert Leboe.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Name Walter Henderson Name Walter Clarence Henderson
Profession Farmer Profession Farmer
Party Progressive Conservative Party Progressive Conservative
2. Text: The production of Tautona mine is 235,000 ounces in 2013.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Title Tautona mine Title Tautona mine
Subtitle Production Subtitle Production
Year 2013 Year 2013
Production (ounces) 235,000 Production 235,00 ounces
3. Text: Elango Nagarajah, also known as “Thaimann Elango", is a Tamil film actor, director, producer and lyricist in the Tamil film industry. He began his career in his early ages as a producer for the Tamil film Anbudan, starred Arun Vijay, Meena, Rambha (actress) in the main was released in the year 2000.
Generated Table:
Predicted Header Prediction - iTBLS Ground Truth Header Ground Truth
Name Elango Nagarajah Name Elango
Occupation actor, director, producer, lyricist Known for Thaimann
Table 6: Difference between the tables generated by the Zero Shot (ZS) and Fine-Tuned (FT) approaches with respect to the Ground Truth on the WikiBio and WikiTableText datasets with additions and deletions represented using red and green.
6 Conclusion
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations addressing three types of tasks – interpretation, modification, and generation. In contrast to prior tabular datasets that are sourced from Wikipedia or financial reports, iTBLS is situated in tabular data obtained from scientific pre-prints on ArXiv. Success on the iTBLS dataset requires understanding both ordinal and cardinal references to cell positions, and understanding implicit references. Additionally, the paper introduces a novel framework that reformulates tabular operations as question-answering. Appropriate questions are created based on the input table and the nature of interaction, and the user request is used as evidence to obtain the answers. The developed approach demonstrates an improvement over a sequence-to-sequence modeling approach on the iTBLS dataset. In addition, the question-answering-based reformulation is evaluated on datasets for the text-to-table task, obtaining up to 13% improvement in terms of exact-match accuracy and 16% improvement in terms of BERTScore compared to the prior state-of-the-art.
Limitations
While iTBLS introduces a dataset for interactive conversations over tabular information, there are some avenues for improvement. In this dataset, modification is modeled as a series of swaps. A more comprehensive sequence of manipulations includes in-place modification of values and modifying a cell’s value based on other cells using both absolute and relative references. While sourcing tabular information from arXiv provides a cost-efficient approach, LLMs are often pre-trained on sources from arXiv. This paper alleviates the issue by sourcing natural language commands from crowdworkers. Future work could look at collecting tabular information from crowdworkers as well. While we present a suite of baseline approaches for iTBLS, there is still headroom between the presented approaches and perfect performance. We identify the closure of this gap as an avenue for future work.
Acknowledgments
This work was supported by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
Appendix A Appendix
A.1 AI Assistance Acknowledgment
We acknowledge the use of GitHub Copilot to assist in code completion.
A.2 Compute
All fine-tuning and inference was run on Nvidia A40 GPUs with 48GB GDDR6 memory. Fine-tuning took  1-2 hours on 8 GPUs in parallel with pytorch distributed data parallel (DDP).
A.3 Dataset Statistics
Statistics of the text-to-table datasets:
Dataset Train Valid Test
E2E 42.1k 4.7k 4.7k
WikiTableText 10k 1.3k 2.0k
WikiBio 582.7k 72.8k 72.7k
Table 7: Statistics of the E2E, WikiTableText, WikiBio, and RotoWire datasets, number of samples across splits
A.4 Dataset Examples – Text to Table
This section details example textual paragraphs and associated tables from the different datasets.
E2E:
The Eagle is a low rated coffee shop near Burger King and the riverside that is family friendly and is less than £20 for Japanese food.
Name The Eagle
Food Japanese
Price range Less than £20
Customer Rating Low
Area Riverside
Family friendly Yes
Near Burger King